### v1

@@ 2. Study Design
In RQ1, we investigated the effectiveness of each studied open-source LLM in unit test generation by designing different prompts. This helps investigate the influence of different prompt designs on this task and also helps determine how to design a prompt to make LLMs achieve the best effectiveness. In RQ2, we compared the studied open-source LLMs in unit test generation based on the best prompt design we explored in RQ1, and employed the state-of-the-art commercial LLM (i.e., GPT-4) and the widely-studied traditional approach (i.e., Evosuite) as reference. This helps understand the current situation of these open-source LLMs in unit test generation. In RQ3, we incorporated several widely-used ICL methods to improve the usage of these studied LLMs respectively, and then revisited the effectiveness of these LLMs in unit test generation. This is helpful to explore whether the state-of-the-art ICL methods in general tasks can help improve the effectiveness of LLMs in the specific task of unit test generation.
The former three RQ3 correspond to the above-mentioned three aspects affecting the effectiveness of LLMs in addressing some tasks, which adopted the widely-used evaluation metrics in existing studies on LLM-based unit test generation (i.e., syntactic validity and test coverage). RQ4 further complements them by considering another key ability (i.e., the defect detection ability) of unit tests. That is, in RQ4, we investigated the effectiveness of these studied LLMs in terms of fault detection with the corresponding automatically generated unit tests.
@@ 2.6. Implementation
We developed our experimental scripts based on PyTorch 2.0.0 (tor, 2023) and transformers 4.34.1 (tra, 2023), and used DeepSpeed (Rasley et al., 2020) and VLLM (Kwon et al., 2023) libraries to speed up our experiments. All the experiments were conducted on four servers with the same configuration, i.e., Ubuntu 18.04 LTS, Intel Xeon Gold 6240C CPU, 512GB RAM, and eight NVIDIA A100 GPUs. All of our code and data are available at our project homepage (hom, 2024).
@@ 2.7. Metrics
Following the existing studies (Yuan et al., 2023b; Xie et al., 2023), we leveraged the three widely-used metrics to evaluate the generated unit tests by each technique in RQ1, RQ2, and RQ3: Compilation Success Rate (CSR), Line Coverage ($Cov_{L}$), and Branch Coverage ($Cov_{B}$). As introduced in Section 2.6, for each focal method, we integrated all the generated unit tests into a test class, and then compiled the test class. Therefore, we calculated CSR by the ratio of the number of test classes compiled successfully over the total number of test classes corresponding to all focal methods. Then, we executed each test class compiled successfully, and collected its achieved line coverage and branch coverage via the code coverage collection tool (i.e., Jacoco (Hoffmann and Ernst, 2021)). For each test class with compilation errors, we recursively remove test methods occurred in compilation error messages from the test class and re-execute the test class until it can be compiled successfully or there is no test method left in the test class. In this way, we try our best to collect as much coverage achieved as possible. Higher CSR, $Cov_{L}$, and $Cov_{B}$ indicate better effectiveness in unit test generation.
@@ 3.2. Effectiveness Comparison
Table 4 shows the effectiveness of each studied open-source LLM with the corresponding best setting mentioned above in terms of CSR, $Cov_{L}$, and $Cov_{B}$. To better understand the effectiveness of open-source LLMs in unit test generation, we also reported the effectiveness of traditional technique Evosuite with its default setting, as well as the closed-source GPT-4 model by incorporating all studied code features in the natural language description style (the most widely-used setting in the literature (Yuan et al., 2023b; Xie et al., 2023; Tang et al., 2023)).
@@
To potentially improve the effectiveness of LLM-based unit test generation, we further analyzed the underlying root cause of their low CSR. Specifically, there are three main types of invalid unit tests generated by these LLMs: unresolved symbol errors, parameter mismatch errors, and abstract instantiation errors. All of them are caused by the hallucination issue (Liu et al., 2024) of LLMs. Unresolved symbol errors are the most common, which account for 30.68% of all invalid unit tests and are caused due to using semantically-coherent but undefined identifiers in unit tests. Parameter mismatch errors are the second most common, accounting for 17.25%, which involves incorrect parameter types or counts in invoked APIs. This is because there are many APIs sharing the same name but different parameter lists due to the overloading feature, aggregating the difficulty in generating correct API invocations for LLMs. Abstract instantiation errors are the third most common, accounting for 10.38%, which involves incorrect instantiation of abstract classes. Specifically, when a method parameter is defined in an abstract class, an instance of the corresponding subclass should be declared since abstract classes cannot be instantiated. However, LLMs often declare an instance for the parameter according to its defined type.
@@ 3.3. Effectiveness of In-Context Learning
Finding 6: CoT is helpful to improve the effectiveness of both DeepSeek-Coder models in unit test generation while it cannot for the three CodeLlama models due to the stronger ability of the former in code comprehension.
@@ 3.4. Defect Detection Ability
We then reported the number of detected defects by each LLM in Row “NDD” in Table 6. We further confirmed the larger-scale LLMs (including GPT-4, PD-34B, and DC-33B) detected more defects than the remaining smaller-scale LLMs. Among these detectable defects, however, the percentage of detected defects is still small, ranging from 29.27% to 60.00%. This indicates the weak defect detection ability of LLM-based unit test generation.
Finding 8: The defect detection ability of LLM-based unit test generation is weak, mainly caused by the low validity of generated unit tests. On average, for 87.13% defects, LLMs do not generate any valid unit tests. Among the remaining defects, just 47.28% defects are exactly detected.
We then analyzed why just a few defects are detected among all detectable ones in order to further understand the challenges for LLM-based unit test generation in defect detection. Through our manual investigation on each valid unit test corresponding to each undetected defect, we summarized the following three reasons. Table 7 presents the number of undetected defects caused by each reason for each LLM.
Table 7. Number of Defects that Valid Unit Tests Failed to Detect Categorized by Three Reasons
@@
Improper Assertionw 1 0 1 0 0 2
(1) Insufficient test coverage refers to that all the generated unit tests by an LLM do not reach the faulty code. From Table 7, there are 3.61%, 1.81%, 6.63%, 6.02%, 5.42% of undetected defects caused by this reason for CL-7B, CL-13B, PD-34B, DC-7B, DC-33B, respectively. (2) Missing specific inputs refers to that there are some unit tests executing the faulty code but the inputs specified in the unit tests cannot trigger the defect. That is, the triggering of the defect requires some specific inputs (such as setting the real part of a complex number to NaN for triggering the defect in Math-53). From Table 7, there are 13.25%, 6.02%, 12.65%, 15.66%, 12.05% , 14.46% of undetected defects caused by this reason for CL-7B, CL-13B, PD-34B, DC-7B, DC-33B, GPT-4, respectively. (3) Improper assertions refer to that the defect-triggering inputs have been specified in some unit tests, but no proper assertions are produced to capture this defect. For example, the defect in Compress-34 lies in missing to flush the input stream after closing a file. However, the unit tests generated by PD-34B do not contain the assertions checking whether the corresponding attribute is null, but just check whether the file is closed. From Table 7, there are 0.60%, 0.60%, 1.20% of undetected defects caused by this reason for CL-7B, PD-34B, GPT-4, respectively.
@@ 6. Related Work
Unit Test Generation. In the literature, a significant body of research has been dedicated to the development of automated unit test generation techniques (Serra et al., 2019; Tufano et al., 2020; Alagarsamy et al., 2023; Yuan et al., 2023b; Dakhel et al., 2024; Xie et al., 2023; Siddiq et al., 2023; Bhatia et al., 2023; Li and Yuan, 2024). They can be divided into two main categories: traditional techniques and DL-based techniques.
@@
Recently, with the huge impact of ChatGPT, researchers explored the effectiveness of LLMs in generating unit tests. Specifically, Zhang et al. (Yuan et al., 2023b) proposed Chattester that iteratively generates unit tests based on conversations with ChatGPT. Lemieux et al. (Lemieux et al., 2023) proposed CODAMOSA that combines evolutionary search with CodeX (Chen et al., 2021) by exploiting code understanding ability of CodeX to overcome the low rate of generating coverage-improving tests with traditional methods. Li et al. (Li et al., 2023c) proposed using the subtle differences between fixed and buggy code versions to guide ChatGPT to generate failure-inducing tests. These LLM-based techniques show significant superiority over earlier DL-based ones. However, they are all based on closed-source LLMs, resulting in a lack of transparency and reproducibility. Particularly, it is still unclear how recent open-source LLMs perform in unit test generation.
Empirical Study on evaluating LLMs in Code-Related Tasks. With the rapid development in LLMs, there are multiple studies investigating the effectiveness of LLMs in code-related tasks (Du et al., 2024; Nam et al., 2024; Sun et al., 2023; Molina and Gorla, 2024; Xia et al., 2022; Niu et al., 2023; Wang et al., 2024; Pan et al., 2024; Li et al., 2024; Liu et al., 2023b; Hou et al., 2023; Mastropaolo et al., 2023; Li et al., 2023a; Guo et al., 2024). In particular, there are some empirical studies on LLM-based unit test generation (Yuan et al., 2023b; Xie et al., 2023; Siddiq et al., 2023; Schäfer et al., 2024; Dakhel et al., 2024). For example, Yuan et al. (Yuan et al., 2023b) investigated the effectiveness of ChatGPT and proposed ChatTester by employing CoT for unit test generation. Xie et al. (Xie et al., 2023) proposed and evaluated ChatUniTest by employing the feedback-based generation process with ChatGPT. Siddiq et al. (Siddiq et al., 2023) investigated the effectiveness of GPT-3.5 and Codex in unit test generation. Schafer et al. (Schäfer et al., 2024) proposed TestPilot, which incorporates API documentation and GPT-3.5 for unit test generation.
@@
Furthermore, there are some empirical studies on evaluating LLMs in addressing other code-related tasks. For example, Du et al. (Du et al., 2024) conducted experiments to investigate the effectiveness of LLMs in class-level code generation. Nam et al. (Nam et al., 2024) compared the code understanding ability of LLMs when using them to help developers understand the project. Sun et al. (Sun et al., 2023) conducted experiments to investigate the effectiveness of ChatGPT in code summarization. Facundo et al. (Molina and Gorla, 2024) summarized three types of oracles and threats that arise from automatically generating oracles using LLMs. Xia et al. (Xia et al., 2022) conducted experiments on LLMs to investigate their effectiveness in automated program repair. Niu et al. (Niu et al., 2023) conducted a comprehensive investigation on the effectiveness of pre-trained models in six code-related tasks. Different from them, our work targets the task of unit test generation.

### v2

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: 39th IEEE/ACM International Conference on Automated Software Engineering ; October 27-November 1, 2024; Sacramento, CA, USAbooktitle: 39th IEEE/ACM International Conference on Automated Software Engineering (ASE ’24), October 27-November 1, 2024, Sacramento, CA, USAdoi: 10.1145/3691620.3695529isbn: 979-8-4007-1248-7/24/10ccs: Software and its engineering Software testing and debugging
@@ 2.1. Research Questions
In RQ1, we investigated the effectiveness of each studied open-source LLM in unit test generation by designing different prompts. This helps investigate the influence of different prompt designs on this task and also helps determine how to design a prompt to make LLMs achieve the best effectiveness. In RQ2, we compared the studied open-source LLMs in unit test generation based on the best prompt design we explored in RQ1, and employed the state-of-the-art commercial LLM (i.e., GPT-4) and the widely-studied traditional approach (i.e., Evosuite) as reference. This helps understand the current situation of these open-source LLMs in unit test generation. In RQ3, we incorporated two widely-used ICL methods to improve the usage of these studied LLMs respectively, and then revisited the effectiveness of these LLMs in unit test generation. This is helpful to explore whether the state-of-the-art ICL methods in general tasks can help improve the effectiveness of LLMs in the specific task of unit test generation.
The former three RQs correspond to the above-mentioned three aspects affecting the effectiveness of LLMs in addressing some tasks, which adopted the widely-used evaluation metrics in existing studies on LLM-based unit test generation (i.e., syntactic validity and test coverage). RQ4 further complements them by considering another key ability (i.e., the defect detection ability) of unit tests. That is, in RQ4, we investigated the effectiveness of these studied LLMs in terms of fault detection with the corresponding automatically generated unit tests.
@@ 2.6. Implementation
We developed our experimental scripts based on PyTorch 2.0.0 (tor, 2023) and transformers 4.34.1 (tra, 2023), and used VLLM (Kwon et al., 2023) libraries to speed up our experiments. All the experiments were conducted on four servers with the same configuration, i.e., Ubuntu 18.04 LTS, Intel Xeon Gold 6240C CPU, 512GB RAM, and eight NVIDIA A100 GPUs. All of our code and data are available at our project homepage (hom, 2024).
@@ 2.7. Metrics
Following the existing studies (Yuan et al., 2023b; Xie et al., 2023), we leveraged the three widely-used metrics to evaluate the generated unit tests by each technique in RQ1, RQ2, and RQ3: Compilation Success Rate (CSR), Line Coverage ($Cov_{L}$), and Branch Coverage ($Cov_{B}$). As introduced in Section 2.6, for each focal method, we integrated all the generated unit tests into a test class, and then compiled the test class. Therefore, we calculated CSR by the ratio of the number of test classes compiled successfully over the total number of test classes corresponding to all focal methods. Then, we executed each test class that was successfully compiled, and collected its achieved line coverage and branch coverage via the code coverage collection tool (i.e., Jacoco (Hoffmann and Ernst, 2021)). For each test class with compilation errors, we recursively remove test methods occurred in compilation error messages from the test class and re-execute the test class until it can be compiled successfully or there is no test method left in the test class. In this way, we try our best to collect as much coverage achieved as possible. Higher CSR, $Cov_{L}$, and $Cov_{B}$ indicate better effectiveness in unit test generation.
@@ 3.2. Effectiveness Comparison
Table 4 shows the effectiveness of each studied open-source LLM with the corresponding best prompt design mentioned above in terms of CSR, $Cov_{L}$, and $Cov_{B}$. To better understand the effectiveness of open-source LLMs in unit test generation, we also reported the effectiveness of traditional technique Evosuite with its default setting, as well as the closed-source GPT-4 model by incorporating all studied code features in the natural language description style (the most widely-used setting in the literature (Yuan et al., 2023b; Xie et al., 2023; Tang et al., 2023)).
@@
To potentially improve the effectiveness of LLM-based unit test generation, we further analyzed the underlying root cause of their low CSR. Specifically, there are three main types of invalid unit tests generated by these LLMs: unresolved symbol errors, parameter mismatch errors, and abstract instantiation errors. All of them are caused by the hallucination issue (Liu et al., 2024; Wei et al., 2023; Agrawal et al., 2023) of LLMs. Unresolved symbol errors are the most common, which account for 30.68% of all invalid unit tests and are caused due to using semantically-coherent but undefined identifiers in unit tests. Parameter mismatch errors are the second most common, accounting for 17.25%, which involves incorrect parameter types or counts in invoked APIs. This is because there are many APIs sharing the same name but different parameter lists due to the overloading feature, aggregating the difficulty in generating correct API invocations for LLMs. Abstract instantiation errors are the third most common, accounting for 10.38%, which involves incorrect instantiation of abstract classes. Specifically, when a method parameter is defined in an abstract class, an instance of the corresponding subclass should be declared since abstract classes cannot be instantiated. However, LLMs often declare an instance for the parameter according to its defined type.
@@ 3.3. Effectiveness of In-Context Learning
Finding 6: CoT is helpful to improve the effectiveness of both DeepSeek-Coder models in unit test generation while it does not benefit the three CodeLlama models due to the stronger ability of the former in code comprehension.
@@ 3.4. Defect Detection Ability
We then reported the number of detected defects by each LLM in the row “NDD” in Table 6. We further confirmed the larger-scale LLMs (including GPT-4, PD-34B, and DC-33B) detected more defects than the remaining smaller-scale LLMs. Among the detected defects, there are two types of defect-triggering reasons: run-time exceptions and assertion errors. For example, DC-7B, PD-34B, DC-33B, and GPT-4 detected 6, 16, 17, and 19 defects triggered by assertion errors, respectively, and 18, 14, 16, and 20 defects triggered by run-time exceptions. On the other hand, all defects detected by CL-7B and CL-13B were triggered by run-time exceptions, showing their relatively weaker ability in generating proper assertions. Despite these findings, the percentage of detected defects is still small, ranging from 29.27% to 60.00%. This indicates the weak defect detection ability of LLM-based unit test generation.
Finding 8: The defect detection ability of LLM-based unit test generation is weak, mainly caused by the low validity of generated unit tests. On average, for 87.13% defects, LLMs do not generate any valid unit tests. Among the remaining defects, only 47.28% defects are exactly detected.
We then analyzed why just a few defects are detected among all detectable ones in order to further understand the challenges for LLM-based unit test generation in defect detection. Specifically, following the PIE theory in software engineering (Voas, 1992), we designed three reasons why these defects were not detected by generated tests. Then, four authors with over four years of Java programming experience manually analyzed and independently labeled each defect. The Cohen’s Kappa score is 0.95, indicating the labeling quality to some degree. Then, all the inconsistencies in labeling were discussed together to reach a consensus. Table 7 presents the number of undetected defects caused by each reason for each LLM.
Table 7. Number of Defects that Valid Unit Tests Failed to Detect Categorized by Three Reasons.
@@
Improper Assertions 1 0 1 0 0 2
(1) Insufficient test coverage refers to that all the generated unit tests by an LLM do not reach the faulty code. From Table 7, there are 3.61%, 1.81%, 6.63%, 6.02%, 5.42% of undetected defects caused by this reason for CL-7B, CL-13B, PD-34B, DC-7B, and DC-33B, respectively. (2) Missing specific inputs refers to that there are some unit tests executing the faulty code but the inputs specified in the unit tests cannot trigger the defect. That is, the triggering of the defect requires some specific inputs (such as setting the real part of a complex number to NaN for triggering the defect in Math-53). From Table 7, there are 13.25%, 6.02%, 12.65%, 15.66%, 12.05% , 14.46% of undetected defects caused by this reason for CL-7B, CL-13B, PD-34B, DC-7B, DC-33B, and GPT-4, respectively. (3) Improper assertions refer to that the defect-triggering inputs have been specified in some unit tests, but no proper assertions are produced to capture this defect. For example, the defect in Compress-34 lies in missing to flush the input stream after closing a file. However, the unit tests generated by PD-34B do not contain the assertions checking whether the corresponding attribute is null, but just check whether the file is closed. From Table 7, there are 0.60%, 0.60%, 1.20% of undetected defects caused by this reason for CL-7B, PD-34B, and GPT-4, respectively.
@@ 6. Related Work
Unit Test Generation. In the literature, a significant body of research has been dedicated to the development of automated unit test generation techniques (Serra et al., 2019; Tufano et al., 2020; Alagarsamy et al., 2023; Yuan et al., 2023b; Dakhel et al., 2024; Xie et al., 2023; Siddiq et al., 2023; Bhatia et al., 2023; Li and Yuan, 2024; Yang et al., 2024). They can be divided into two main categories: traditional techniques and DL-based techniques.
@@
Recently, with the huge impact of ChatGPT, researchers explored the effectiveness of LLMs in generating unit tests. Specifically, Yuan et al. (Yuan et al., 2023b) proposed ChatTester that iteratively generates unit tests based on conversations with ChatGPT. Lemieux et al. (Lemieux et al., 2023) proposed CODAMOSA that combines evolutionary search with CodeX (Chen et al., 2021) by exploiting code understanding ability of CodeX to overcome the low rate of generating coverage-improving tests with traditional methods. Li et al. (Li et al., 2023c) proposed using the subtle differences between fixed and buggy code versions to guide ChatGPT to generate failure-inducing tests. Schafer et al. (Schäfer et al., 2024) proposed TestPilot, which incorporates API documentation and GPT-3.5 for unit test generation. These LLM-based techniques show significant superiority over earlier DL-based ones. However, they are all based on closed-source LLMs, resulting in a lack of transparency and reproducibility. Recently, Yang et al. (Yang et al., 2024) proposed TELPA, which aims to improve the effectiveness of LLMs in unit test generation by integrating them with bidirectional method invocation analysis and counterexample sampling. However, their evaluation was limited to PD-34B, and it remains unclear how recent open-source LLMs perform in unit test generation.
Empirical Study on evaluating LLMs in Code-Related Tasks. With the rapid development in LLMs, there are multiple studies investigating the effectiveness of LLMs in code-related tasks (Du et al., 2024; Nam et al., 2024; Sun et al., 2023a; Molina and Gorla, 2024; Xia et al., 2022; Niu et al., 2023; Wang et al., 2024; Pan et al., 2024; Li et al., 2024; Liu et al., 2023b; Hou et al., 2023; Mastropaolo et al., 2023; Li et al., 2023a; Guo et al., 2024; Chen et al., [n. d.]). In particular, there are some empirical studies on LLM-based unit test generation (Yuan et al., 2023b; Xie et al., 2023; Siddiq et al., 2023; Schäfer et al., 2024; Dakhel et al., 2024). For example, Yuan et al. (Yuan et al., 2023b) investigated the effectiveness of ChatGPT in terms of the correctness, sufficiency, readability and usability of generated unit tests in unit test generation. Siddiq et al. (Siddiq et al., 2023) investigated the effectiveness of GPT-3.5 and Codex in unit test generation. Schafer et al. (Schäfer et al., 2024) investigated the effectiveness of GPT-3.5 in generating unit tests for Javascript language..
@@
Furthermore, there are some empirical studies on evaluating LLMs in addressing other code-related tasks. For example, Du et al. (Du et al., 2024) conducted experiments to investigate the effectiveness of LLMs in class-level code generation. Nam et al. (Nam et al., 2024) compared the code understanding ability of LLMs when using them to help developers understand the project. Sun et al. (Sun et al., 2023a) conducted experiments to investigate the effectiveness of ChatGPT in code summarization. Facundo et al. (Molina and Gorla, 2024) summarized three types of oracles and threats that arise from automatically generating oracles using LLMs. Researchers (Xia et al., 2022; Jiang et al., 2024) conducted experiments on LLMs to investigate their effectiveness in automated program repair. Niu et al. (Niu et al., 2023) conducted a comprehensive investigation on the effectiveness of pre-trained models in six code-related tasks. Different from them, our work targets the task of unit test generation.
Test Oracle Generation Due to the importance of test oracles, numerous research efforts have been dedicated to automatically generating test oracles for test cases (Dinella et al., 2022; Watson et al., 2020; Yu et al., 2022; Sun et al., 2023b). For example, Watson et al. (Watson et al., 2020) proposed ATLAS, which adopted a bidirectional RNN for test assertion generation. Dinella et al. (Dinella et al., 2022) introduced TOGA to infer exceptional and assertion test oracles based on the context of the focal method. Yu et al. (Yu et al., 2022) first proposed an IR-based method for test assertion generation, and Sun et al. (Sun et al., 2023b) further developed EDITAS to improve IR-based test assertion generation by considering the differences between the retrieved and target “focal-test” method pairs. In our study, we use LLMs to generate complete unit tests, including both test prefixes and test oracles. When measuring the effectiveness of generated unit tests (e.g., validity and defect-detection effectiveness), we also considered both test prefixes and oracles.
@@ 7. Conclusion
Acknowledgements.
We gratefully acknowledge the financial support provided by the National Natural Science Foundation of China (Grant Nos. 62322208, 62202040, 62232001, 12411530122) and CCF-Huawei Populus Grove Fund which made this research possible. We would like to express our sincere appreciation to the anonymous reviewers for their constructive suggestions, which have greatly improved the quality of this paper.